OpenStudy (anonymous):

Using complete sentences, explain how to find the zeros of the function f(x) = 2x3 – 9x + 3.

OpenStudy (anonymous):

You need to equal F(x) to 0 (zero). So, f(x)=9-9x <=> f(x)=9(1-x) F(x) = 0 <=> 9(1-x)=0 <=> 9=0 (false) V 1-x=0 <=> x=1